1 2 3 4 5 6 7 8 9  1 2 3 4 5 6 7 8 9 1 2 3 4 5 6 7 8 9 1 2 3 4 5 6 7 8 9

The frame on this page should be 24,6501778cm×15,9424103cm = 9,7047944′′ × 6,2765395′′. When printed on A4 paper, the frame should be one inch from each edge of the paper.

The rulers at the four edges will indicate how much of the page is useable. The large ticks are .1′′ apart.